Sydney purchased a lottery ticket and won $5,000.00. She invested part of her winnings in a safe account that paid 3% interest per year and the rest of the winnings in a risky venture that ended up paying 7% at the end of the year. At the end of the year she had a profit of $270. How much did she invest in the risky venture?
I identified what my variable is r = risky venture
I believe I have to use I=P*R*T
I=5000*6*1 and
I=5000*7*1

Answer by mananth(16949)

risky venture amount r
interest from it = 0.07r
Balance at 3%
5000-r @ 3%
Interest = 0.03(5000-r)
total interest = 270
0.07r+0.03(5000-r)=270
0.07r+150-0.03r=270
0.04r= 270-150
0.04r=120
r=120/0.04
r=3000